1. Evaluate "\intop"x2(1 + 2x3)3dx.


2. Evaluate "\intop"xe7x dx.

3. Find the volume of the solid of revolution when the curve y = 1 + x2 is revolved around the x-axis on [−2, 2].


1
Expert's answer
2021-12-17T12:33:02-0500

(1) "I=\int x^{2}\left(1+2 x^{3}\right)^{3} d x"

By substitution method, let "u=1+2 x^{3}"

Such that "d u=6 x^{2} d x"

"\begin{aligned}

&I=\int x^{2}\left(1+2 x^{3}\right)^{3} d x=I=\int x^{2} u^{3} \frac{d u}{6 x^{2}} \\

&I=\frac{1}{4} \int u^{3} d u=\frac{u^{4}}{24}+C \\

&I=\frac{\left(1+2 x^{3}\right)^{4}}{24}+C

\end{aligned}"

(2) "I=\int x e^{7 x} d x"

Using integration by part

"\begin{aligned}

&u=x \quad, \quad d v=e^{7 x} d x \\

&d u=1, \quad v=\frac{e^{7 x}}{7} \\

&\int u d v=u v-\int v d u \\

&I=\frac{x e^{7 x}}{7}-\frac{1}{7} \int e^{7 x} d x

\end{aligned}"


"I=\frac{x e^{7 x}}{7}-\frac{e^{7 x}}{49}+C"

(3) "y=1+x^{2}"

Using method of disk

"\begin{aligned}

&V=\int_{-2}^{2} \pi y^{2} d x=\int_{-2}^{2} \pi\left(1+x^{2}\right)^{2} d x \\

&V=\int_{-2}^{2} \pi\left(1+2 x^{2}+x^{4}\right) d x \\

&V=\frac{412 \pi}{15}

\end{aligned}"